OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 2000 centimeters/minute while the area of the triangle is increasing at a rate of 2000 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 8000 centimeters and the area is 96000 square centimeters?

OpenStudy (ageta):

A = bh/2 = 99000 = bh/2 h = 10000 198000 = 10000b b = 19.8 A = bh/2 dA/dt = (bdh/dt + hdb/dt)/2 A' = (bh' + hb')/2 h' = 1500 cm/min A' = 1000 cm²/min 1000 = (1500b + b'h)/2 2000 = 1500b + b'h b' = (2000 - 1500b)/h b' = (2000 - 1500*19.8)/10000 b' = db/dt = -2.77 cm/min n